Circular No. 6120

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

SUPENOVA 1994ai IN NGC 908 A. Williams and R. Martin report their discovery at the Perth Ob-

servatory of an apparent supernova located ∼ 24′′ east and 4′′ north of the center of NGC 908 (α = 2^{h}20^{m}46^{s}, δ = −21^{o}27′.6, equinox 1950.0). The object was found on an image taken on Dec. 20.52 UT in the course of the Perth Astronomy Research Group’s automated supernova search. Estimated unfiltered magnitudes: Dec. 8, [19; 16, 17–18; 20.52, 16–17.

S. Benetti, European Southern Observatory, reports that inspection of a preliminarily-reduced CCD spectrum (range 400-800 nm, resolution 2.6 nm), obtained on Dec. 24.11 UT with Danish 1.54-m telescope (+ DFOSC) at La Silla, confirms that SN 1994ai is a supernova of type Ic around maximum. Lines (with P-Cyg profile) of Fe II, Na I D, Si II 635.5-nm, and O I 777.4-nm are superimposed on a red continuum. The expansion velocity deduced from the minimum of the Na I D, Si II, and O I lines, corrected by the recession velocity of 1700 km/s of the parent galaxy (measured from the Na I D interstellar lines), is ∼ 8000 km/s.

SUPERNOVA 1993J IN NGC 3031 H.-U. Zimmermann, B. Aschenbach, G. Hasinger, W. Pietsch, P. Pre-

dehl, and J. Trümper, Max-Planck-Institut für Extraterrestrische Physik; W. Lewin, Massachusetts Institute of Technology; J. van Paradijs and E. Magnier, University of Amsterdam; G. Fabbiano, Center for Astrophysics; L. Lubin, Princeton University; and R. Petre, Goddard Space Flight Center, report: “SN 1993J was again observed with ROSAT between Oct. 19 and 21; during a 20 000-s observation with the HRI detector (which has no spectral resolution), a mean rate of 0.0072 count/s was recorded. Under the assumption that the spectral shape in the ROSAT band (0.1–2.4 keV) has not drastically changed since April 1994 (cf. IAUC 6014), the soft-x-ray flux has remained roughly constant since then. This may indicate that the shock wave has now entered regions with higher matter densities.”

EX HYDRAE A. Jones, Nelson, New Zealand, reports that this cataclysmic variable

is in outburst, as indicated by his visual magnitude estimates: Dec. 24.58 UT, 13:; 25.60, 9.6; 25.64, 9.4.

1994 December 26 Daniel W. E. Green